OpenStudy (anonymous):

a password of length 4 that uses exactly 2 different digits is to be created. How many different passwords can there be? here, 0001 is a valid password, while 3333 and 0123 are not valid passwords because they do not use exactly 2 different digits. @cwrw238 @amistre64

OpenStudy (amistre64):

10*10*8*8 maybe?

OpenStudy (anonymous):

please do explain... i thought of this considering two cases , one wid 3 same and 1 same digit ,, and the other with 2 same and 2 same digits... And the answer is to be a max of 3 digits....

OpenStudy (amistre64):

1000 0100 1100 1010 1001 0110 0101 0011 1110 1101 1011 0111 that seems to be the set of any 2 digit combination; and to avoid using the same digit twice, there are 9 of those sets which most likely doubles the lot since the set of say 2,9 is the same as 9,2 just taking a stab at it

OpenStudy (amistre64):

hmm, or is it 10 P 2 sets; how many ways are there to pick 2 numbers from a set of 10 ?

OpenStudy (amistre64):

1000 0100 0100 0001 <-- forgot to include the last 2 :) 1100 1010 1001 0110 0101 0011 1110 1101 1011 0111

OpenStudy (anonymous):

ya there wud be 10 P 2 i.e. 10 C2 * 2! ways for picking 2 out of 10 numbers...

OpenStudy (anonymous):

\[\left(\begin{matrix}10 \\ 2\end{matrix}\right) * (\left(\begin{matrix}4 \\ 1\end{matrix}\right) + \left(\begin{matrix}4 \\ 2\end{matrix}\right) + \left(\begin{matrix}4 \\ 3\end{matrix}\right)) = 630\]

OpenStudy (amistre64):

i forget the setup, but spose you had 10 people, how many different groups of 2 people could you create?

OpenStudy (anonymous):

Hey cnknd , can you explain ur method please???

OpenStudy (amistre64):

01 02 03 04 05 06 07 08 09 12 13 14 15 16 17 18 19 23 24 25 26 27 28 29 34 35 36 37 38 39 45 46 47 48 49 56 57 58 59 67 68 69 78 79 89 45 different setups

OpenStudy (anonymous):

ok 10 choose 2 gives u 45 combinations of 2 digits, and the other factor is what @amistre64 wrote in a previous post: 1000 0100 0100 0001 1100 1010 1001 0110 0101 0011 1110 1101 1011 0111

OpenStudy (amistre64):

and 14 different setups for each group; 45*14 ... brute method ;)

OpenStudy (anonymous):

O.k . thank u both, I am trying to understand the method... hoping will not face problem...

OpenStudy (amistre64):

10 choose 2 ultimate setups within each setup 4 choose 1 (0001 0010 0100 1000) + 4 choose 2 (0011 0101 1001 ... ) + 4 choose 3 (1110 1101 1011 0111) so as cnk put it 10C2 times (4C1+4C2+4C3)

OpenStudy (anonymous):

Yes thank u so much...